identify.library

A library that converts cryptical IDs to a human-readable form Version 3.1, 2 August 1996

Richard Körber

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 by Richard Körber Please read the Copyright chapter!

1 Introduction

The identify.library has a rather long history.

It all began when I got mad about all the System info programs only showing the expansion boards as a number. So I wrote expname.library, which may already be known to you.

Later, expname could also identify the system’s parameters (even if it was wrong sometimes ;) ). I still have new ideas for this library. The only trouble is that the name does not fit any more.

Well, now you have the second generation of the expname.library, called identify.library. It has been completely rewritten and expanded. The conceptional flaws of the expname.library has been removed, and a lot of newbies have been added.

The library has currently these features:

• Converting manufacturer and product id into manufacturer name, product name, and product class.

• Converting the cryptical Amiga Guru codes into readable descriptions.

• Analyzing the system’s configuration (system type, processor, memory, graphics and so on...).

Even if the library appears to be quite ready — it is still far away!

I need your help! Please send manufacturer and product names and IDs, your system’s configuration, suggestions, bug reports and so on; simply all you want to see implemented in the library.

2 Copyright

Please read this part carefully.

COPYRIGHT This software is copyrighted by Richard Körber. That means that you are NOT AL-LOWED to modify the program and documentation in any way. Especially you MUST NOT REMOVE the documentation, include files or this text file. You are NOT allowed to use this software or any part of it for any other purpose than that one mentioned in its documentation, this also includes any fonts, images or samples. You are NOT allowed to decompile any part of the packet. This means the database in special! You need a written permission from the author if you want to extract parts of the documentations or include files for own documentations or include files. The ONLY exception is this copyright note, which may be used freely.

FREEWARE This packet is FREEWARE. You are allowed to use it without any registration fee. Note that you only have the right to use this software, but no rights on the software itself.

DISTRIBUTION This package is freely distributable. That means you are allowed to redistribute this package as long as you follow these points:

• Any re-distribution has to include all files in this archive, without any modifications. You are NOT allowed to add any files to the archive. Exception: if you use the identify.library only, you do not need to distribute all related files if you write a short note into your program’s documentation.

• This package may be freely distributed via BBSs, InterNet/UseNet, software libraries such as Fred Fish’s and Aminet CD-ROM, and other similar electronic channels.

• Disk magazines and services that charge extra for file transfers may NOT distribute it without written permission by the developer!

• You are allowed to use the expname.library even in a commercial product without any written permission of me. Please note in your program’s documentation that you’re using identify.library by Richard Körber.

DISCLAIMER By using this product, you accept the FULL responsibility for any damage or loss that might occur through its use or the inability to use it. The developer of the software can NOT be held responsible.

CONTENTS OF THE PACKAGE

The Identify package is only complete if these files are present:

Identify/catalogs/deutsch/Identify.catalog
Identify/catalogs/Identify.cd
Identify/identify.doc
Identify/identify.doc.info
Identify/Identify-D.dvi
Identify/Identify-D.dvi.info
Identify/Identify-D.guide
Identify/Identify-D.guide.info
Identify/Identify-E.dvi
Identify/Identify-E.dvi.info
Identify/Identify-E.guide
Identify/Identify-E.guide.info
Identify/include/identify_lib.i
Identify/include/fd/identify_lib.fd
Identify/include/inline/identify.h
Identify/include/libraries/identify.h
Identify/include/libraries/identify.i
Identify/include/pragmas/identify_pragmas.h
Identify/include/proto/identify_protos.h
Identify/libs/expname.library
Identify/libs/identify.library
Identify/Guru
Identify/ListExp
Identify.info
FILE_ID.DIZ

TRADEMARKS The copyright and trademarks of all manufacturers and products mentioned are held by their owners.

If you do not agree with these copyright notes, you have to delete the Identify packet and all related files immediately!

3 My Address

Please send all kind of bug-reports, board descriptions, missing graphic OS, keyfiles, flames and so on to one of the following e-mail addresses:

INTERNET

I’m reachable trough Internet via these E-Mail addresses:

       rkoerber@tfh.dssd.sub.org
       shred@eratosth.dssd.sub.org

a2h4ko@bg.bib.de (until September 1996 only)

You can also visit my home page (until September 1996 only). There you’ll find the latest version of all my software.

http://www.bg.bib.de/~a2h4ko

SNAIL MAIL

You can send me a snail mail letter, too. My address is:

         Richard Körber
         Hornstraße 20
         51465 Bergisch Gladbach
         Germany

Please enclose a "1 DM" stamp if you live in Germany and want to get a reply.

SUPPORT-MAILBOX You can download the latest version from my support BBS:

                                      ____
____________/   /                  (TM)
\_______/___________________________/   /    ____________________
|  _/  /   /_/   /  _____/___/___/   /    /___/___/  _____/
|     /___/  __/__/   /   /   /   /____/   /   /   /  __/__
:  |\   \/ /___/_______/___/___/___/________/___/___/___/_______/
.__| \   \
    \___\                A FUNHOUSE COMMUNICATIONS DIVISON
Port 01: 02204 - 61630  19.200    Port 02: 02204 / 61499  19.200
Port 03: 02204 - 23856  28.800    Port 04: 02204 / 61032  28.800
Port 05: 02204 - 61033  28.800    Port 06: 02204 / 61034  14.400
Port 07: 02204 - 61035  16.800    Port 08: 02204 / POINT  14.400
Port 09: 02204 - 962121 64.000    Port 10: 02204 / 962122 64.000

4 ListExp

ListExp is a small tool that was already included in the expname.library package.

It describes the hardware components of the system. At first, it will write some hardware parameters (like CPU, Memory). After that, all expansion boards will be listed up (hopefully by name).

By the option FULL, all currently used commodities are listed up, too. The output is then useable for bug reports and similar.

A further option allows to get the manufacturer name, product name and class from manufacturer ID and product ID. Use the options MANUFID and PRODID to do so.

Example:

      > ListExp MANUFID=514 PRODID=9
      -- ListExp V3.0 -- by Richard Koerber
       Manufacturer: Commodore West Chester
       Product:      A2060
       Class:        ArcNet

If some entries in the expansion list begin with a ‘#’ instead of a name, you should send a copy of the ListExp output and the description of the missing board (manufacturer name, product name, product class) to me. :-)

It is also not possible to check out all combinations of Amiga expansions. If your hardware is not properly recognized, please write me.

Please note: Some manufacturers, especially GVP and Phase 5, assigned the same ID to different boards. identify.library is prepared and tries to differ these boards, anyhow it can’t in any case decide right.

5 Guru

Guru translates the cryptical alert code into a human-readable form.

The only parameter is the alert code (all eight digits without any symbols before or behind). An example:

      > guru 81000005
      -- Guru V1.0 -- by Richard Koerber
       Type:       Deadend
       Subsystem:  exec.library
       General:    General fault
       Specified:  Corrupt memory list detected in FreeMem

Please understand the result like that:

Type What kind of alert is it? It may be a deadend alert, or a recovery alert.

Subsystem

Which system caused the alert?

General To what general class does this alert belong? In this case, it is a general fault which must be specified later.

Specified

What was the cause? This example shows that the memory list has been corrupted by a bad-written program.

Appendix A Known Bugs

Identify is written on a Amiga 4000/o3o/882/MMU, Kick 3.0, 14 MB RAM, FastLane 1GB HD, MaestroPro, Toccata, BSC MultiFaceII, CyberVision 4MB.

THESE BUGS ARE CURRENTLY KNOWN:

• ‘Phase 5’ and ‘GVP’ have sometimes assigned one ID to two or more boards. In this case, identify.library may return a name that does not fit to the hardware.

• Currently, an 68030 will only be recognized with MMU.

• I couldn’t test the Amiga analysis completely, because of missing Amiga. If your Amiga model is not properly recognized, please send me a copy of your ListExp output.

Appendix B History

V 3.0 • Bases on the expname.library, but completely rewritten.

• Now convert alert codes, too.

• Wrote GURU.

V 3.1 • 2 boards added

• Bugfix: always recognized an Amiga 3000 on ECS machines with OS3.1.

• Bugfix: recognized CyberGraphX even when PictDT was installed only.

• Improved memory rounding

Appendix C Credits

I want to thank especially these fellows:

Sven Arke for his support with his BBS.

Andreas Gelhausen for Scout, his board list and his ideas.

Thomas Kessler for ShowConfigEnh and his many ideas.

Geert Uytterhoeven for his board list and the description of the GVP codes.

Colin Thompson for his field testings and his useful hints.

Frank Wille for his great PhxAss assembler, and his great hints!

Kössi for his many boards, his help for all my TexInfo problems, and his debug skills.

…and for their contributions (sorted alphabetically):

’2bros’, Daniel Adolfsson, ’ALV’, Thomas Andersson, Jörn Asmussen, Matthias Bethke, Paul Braithwaite, Dave Clarke, Gary Coleman, Norbert Dimpfl, Kevin Fairhurst, Ulrich Falke, Alexander Fichtner, Ramiro Garcia, Andreas Gelhausen, Fred Hamilton, Michael Hartmann, Dirk Hebisch, Matthias Heilmann, Ian P.Heitmans, Gene Heskett, Patrick Hess, Bengt Jensie, Thomas Kessler, Bernd Kösling, James Kückmann, Rask Lambertsen, Jens Langner, Lutz Legero, Petter Lindquist, Mika Lundell, Gunther Mannigel, Peter Mattsson, L Mac Mullan, David Oakes, Jürgen Ofner, Jakob Ölund, Chris Painter, Markus Pietz, Heiko Polig, Michael Reichenbach, Sascha Reissner, Jan Rembser, Kai Rode, Andre Schenk, Thomas Schürger, Simon Shead, Karsten Soeth, Nicholas Stallard, Johan Sundstrom, Adam Szymczak, Patrick Thato, Colin Thompson, Henrik Tikanvaara, Geert Uytterhoeven, Jarkko Vatjus-Anttila, Milco Veljanoski, Federico Villata, Christian Wasner, Ralph Wermke, Henrik Wetterstrom, Martin Wietfeld, Frank Wille, Alexander Wild, Beno Zidaric, Rolf Zuercher

                 //
             \\ //   -- Amiga - The Computer for Creative Mind --

Concept Index 10

Concept Index

A

Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Alert descriptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

B

Bugs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

C

Commercial Product . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Contents of the package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Copyright note. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

D

Disclaimer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Distribution via electronic channels. . . . . . . . . . . . . . . . . . 2

E

E-Mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

F

FreeWare . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

G

Guru . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

H

History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Homepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

I

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

L

ListExp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

S

Snail Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Support-Mailbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

T

Trademarks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

U

Unknown Boards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 My Address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4 ListExp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5 Guru . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Appendix A Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Appendix B History. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Appendix C Credits. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10